# Pixel-Wise Recognition for Holistic Surgical Scene Understanding

### v2

@@ 3 Large Needle Driver (LND) 6 Clip Applier (CA)
We present the Holistic and Multi-Granular Surgical Scene Understanding of Prostatectomies (GraSP) benchmark. An Endoscopic Vision dataset comprised of 32 hours of Robot-Assisted Radical Prostatectomies annotated with a hierarchy of multiple long and short-term tasks, including the recognition of surgical phases and steps through time, the instance segmentation of surgical instruments present in a frame, and the detection of the atomic visual actions performed by those instruments. In this section, we explain the details of the dataset creation process, including data collection, annotation, and curation, and we analyze the main statistics of our dataset.
@@ 3.1 Data Collection
The data acquisition process involved recording the surgical procedures of patients undergoing Robotic-Assisted Laparoscopic Radical Prostatectomy at Fundación Santa Fé de Bogotá, with explicit consent obtained for using these visual materials from the respective patients. We did all the data collection protocols under the approval of the ethics committees of all institutions directly involved in this work.
The surgical team utilized an endoscopic camera to capture recordings, generating a 2D image signal of the left view of the endoscope, which they subsequently extracted through an HDMI cable connected to a KARL STORZ (AIDA mini WD 100) high-definition video capture device. This device facilitated real-time storage of procedural information, providing imagery at 720 dots per inch (dpi) and ensuring robust data storage onto an external hard disk. We anonymized all the data collected according to our ethical protocols.
Three highly experienced surgeons performed the surgeries using a da Vinci Si 3000 robot manufactured by Intuitive Inc. (Sunnyvale, California). The equipment included the robot assembled with a high-definition video camera, providing independent viewing for each eye. The team placed the Monopolar Energy Scissors (Monopolar Curved Scissors) in port #1 (right hand). At the same time, they located the Bipolar Energy Forceps (Bipolar Forceps), capable of being fenestrated or configured as Maryland Forceps, in port #2. The third port accommodated a Prograsp Retraction Forceps (Prograsp Forceps).
@@
For the reconstructive stages of the surgery, surgeons removed the previous instruments located in ports #1 and #2 and positioned two robotic Needle Holders (Large Needle Drivers). Additionally, the surgical setup designated two ports for the bedside assistant. One port housed a Laparoscopic Suction cannula (Suction Instrument), and the other contained either a Laparoscopic Intestinal Retraction Forceps, a Laparoscopic Needle Holder, Laparoscopic Scissors designed explicitly for cutting suture material (Laparoscopic Instruments), or a Clip Applier with Plastic Hemostatic clips of the HEM-O-LOCK type.
3.2 Annotation Process
We classify phases and steps as long-term reasoning tasks, following the temporal breakdown outlined in Huaulmé et al. (2021). Phases represent the fundamental segments of surgical procedures, including key stages like preparation, organ manipulation, and anatomical reconstruction. Meanwhile, steps provide a more detailed breakdown within these phases, focusing on specific surgical activities and maneuvers, such as dissecting anatomical landmarks, making incisions, pulling sutures, and packing dissected structures. In the following subsection, we explain the process of defining phases and step categories.
Conversely, our approach for localized atomic action recognition and surgical instrument segmentation focuses on short-term reasoning tasks. On the one hand, this facet includes identifying localized atomic actions within RARP procedures based on the atomic action definition of AVA (Gu et al. (2018)). Our medical experts stated a set of surgical atomic actions based on the most detailed and singular actions that could be performed in surgery by instruments, such as pull, push, and hold, among others. Table 2 displays the final defined set of atomic action categories on the right side. On the other hand, surgical instrument segmentation involves precisely delineating and identifying surgical tools within the scene. We defined Surgical Instrument classes based on the da Vinci Robot user manual and catalog (Intuitive Inc. (2014, 2023)) and the instrument setup used during data collection (Section 3.1). We showcase the final set of surgical instrument categories in Table 2. By categorizing these aspects as short-term reasoning tasks, our dataset represents both the general and detailed facets of the RARP procedure, enabling a multi-level holistic understanding of surgical scenes in varying degrees of granularity. Due to the time-consuming nature of short-term task annotations and considering our resource constraints, we only annotated a subset of keyframes uniformly sampled through time with a fixed temporal stride, as explained in upcoming Subsections 3.2.2 and 3.2.3.
3.2.1 Phase and Step Tasks Annotation
@@
The delineation of phases, steps, and their corresponding nomenclature adhered closely to the procedural framework of radical prostatectomy coupled with bilateral extended lymphadenectomy. This categorization gained validation through the comprehensive insights outlined in the 12th edition of Campbell-Walsh Urology (Partin et al. (2020)). Within this framework, defining surgical phases involved recognizing distinctive stages within the surgical process, each occurring singularly and without repetition. In contrast, surgical steps refer to specific acts undertaken during the surgical process, often occurring multiple times. The aggregation of multiple steps formed a phase and the determination of the number of phases correlated with the sequential execution of surgical steps.
@@
We then established the sequencing and hierarchy of phases based on standardized steps intrinsic to Robot-Assisted Radical Prostatectomy, a procedural protocol routinely practiced at Fundación Santa Fé de Bogotá. Subsequently, our surgical team reviewed the initial version of the phases and steps categories and hierarchies, involving scrutiny by two experienced urologists with specialized proficiency in Robotic-Assisted Surgery. This collaborative evaluation facilitated the refinement of the defined phases and steps, ensuring alignment with the intricate aspects of the surgical technique. The final set of phases and steps categories is portrayed in Table 2, and the hierarchical structure of phases and steps is visually represented in Figure 2 in the Supplementary Material.
@@
After establishing the labels and hierarchies of the long-term tasks, we conducted an annotation procedure that involved a team of two urologist residents specializing in transperitoneal radical prostatectomy surgeries. The annotation team independently annotated each surgical video, defining specific timestamps for the start and end times of phases and step intervals down to the hour, minute, and second, resulting in two annotations per video. In the case of the eight previously annotated surgeries, these were annotated by the team once more, resulting in two sets of annotations per surgery as well. When discrepancies arose between the annotations, the resident annotators collectively reviewed and reconciled differences to establish a consensus regarding the definitive delineation of phases and steps.
@@ 3.2.2 Surgical Instrument Segmentation
We undertook an annotation process for surgical instrument segmentation that prioritizes dataset consistency, following a structured sequence of steps: (1) frame selection, (2) annotation of surgical instrument masks in the selected frames, (3) and a methodical and iterative validation phase aimed at rectifying errors.
Firstly, we selected frames at 35-second intervals within the surgical videos. This selection aimed to maintain a balanced, uniform representation across categories while considering computational and annotation resource constraints. Secondly, we trained a team of 17 annotators based on an expert-designed and highly specific protocol to maintain annotation quality and consistency across surgical instrument annotations. Employing both Label-Studio (Tkachenko et al. (2020-2022)) and Toronto Annotation Suite (Kar et al. (2021)) open-source annotation platforms, annotators delineated instrument masks and labeled them with one of the predefined instrument sub-type categories. Annotators defined instrument regions either by refining predicted masks semi-automatedly or directly without initial mask predictions. We obtained initial surgical instrument mask predictions using MATIS (Ayobi et al. (2023)), a previous work of ours trained in Endovis 2017 (Allan et al. (2019)) and Endovis 2018 (Allan et al. (2020)). For surgeries in our dataset that already had bounding boxes established for the Instrument Detection task in the earlier version (Valderrama et al. (2022)), we provided these boxes to the annotation team. Their tasks were to refine these bounding boxes into segmentation masks, verify the adequate classification of instruments, and look for possibly missing hidden instances. It is noteworthy to add that we ensured consistent annotation conditions among all annotators to mitigate potential biases in the dataset. Similarly, as an aid for instrument instance identification, all annotators had access to 3-second clips centered on each assigned keyframe along with all the surgical videos corresponding to their assigned keyframes.
Thirdly, we followed a three-staged validation phase. In the first stage, four experts in surgical data science reviewed all segmentation masks to identify any errors, such as missing instrument masks, misclassifications, or delineation errors. These experts accurately described these errors to send these instances back to the annotation team for corrections. In the second stage, the annotation team re-annotated instances with identified errors to ensure rectification. Finally, as a third stage following the re-annotation process, the surgical data science experts conducted a second review of the entire dataset, correcting the remaining annotation errors themselves and ensuring further validation. It is important to note that despite utilizing predefined masks in one of the platforms, the experts in surgical data science ascertained that its presence did not impact the quality of annotations.
3.2.3 Atomic Action Recognition
The annotation process for atomic action recognition involved the team of four experts in surgical data science trained explicitly for this task with the guidance of the surgeons team, following a detailed annotation protocol designed by domain experts. The annotation process consisted of three main stages: (1) Estimation of instrument location with bounding boxes and labeling of atomic actions, (2) Spatial refinement and validation by retrieving bounding boxes from the previously precisely annotated segmentation masks, and (3) Annotation consensus and rectification between independent annotators and guidance of experienced surgeons.
We underwent this process using the open-source Toronto Annotation Suite (Kar et al. (2021)) platform to annotate the same previously selected frames for surgical instrument segmentation. Since atomic action labeling explicitly requires temporal context, annotators defined atomic actions based on the 3-second clips centered on the frame of interest as initially done in AVA (Gu et al. (2018)). However, annotators could also access all of the surgical footage when necessary. Hence, the annotators first localized instruments with bounding boxes regardless of the instrument segmentation annotation process. After defining surgical instruments’ spatial locations with bounding boxes, annotators identified and labeled combinations of maximum 3 atomic actions, treating them as discrete states that instruments could simultaneously undertake based on the predefined set of atomic action categories (Table 2).
Subsequently, we spatially aligned the bounding boxes generated from the atomic action labeling with those obtained from instrument segmentation mask annotations. Hence, we kept the refined bounding boxes derived from segmentation masks to locate atomic actions. This simultaneous annotation procedure allowed a further validation of instrument localization annotations by comparing the outcomes between separate groups of annotators. The surgical data science experts carefully revised cases with considerable discrepancies between the annotated instrument segments and instrument bounding boxes for atomic actions.
The labeling of atomic actions for each instrument was performed in parallel by three different annotators. In the case of the eight previously annotated surgeries, the team re-annotated these two times more, also resulting in three sets of annotations per instrument. Consequently, each instrument in the dataset had three sets of atomic action combinations. To reduce biases and ensure robustness in our annotations, we retained individual atomic action categories labeled by at least two annotators (present in at least two of the three sets) as the final atomic action annotations. For instrument instances where the intersection of at least two sets over the union of the three sets was lower than 0.3, the team collectively reviewed and deliberated upon them to establish a final set of atomic actions.
Additionally, we constructed theoretical co-occurrence matrices under the guidance of our experienced surgeons to establish valid presence relationships among the different atomic actions and between actions and instruments (Figure 11 of the Supplementary Material). These matrices ensured logical consistency in actions performed simultaneously by an instrument and verified correlations between specific actions executed by their annotated instrument type. We applied these matrixes to validate the annotated combinations of atomic actions, ensuring compatibility within sets of actions. In cases where conflicting atomic actions arose, or actions were incompatible with the corresponding instrument, the team collectively reviewed and deliberated to establish consensus for re-annotation.
@@ 3.2.3 Atomic Action Recognition
Table 3: Distribution of the number of frames on each defined dataset split. The Table presents the duration in seconds and the number of frames sampled at 1fps of each collected Case in our dataset. Each Case corresponds to an entire surgery performed. We also show our predefined dataset partition into Fold 1 and Fold 2 for two-fold cross-validation and a Test set for final model testing.
Set Case No. Duration (h) Frames (1fps)
Fold 1 001 3.05 10972
004 2.36 8490
014 2.77 9973
015 2.48 8969
Fold 2 002 2.62 9443
003 1.63 5882
007 2.23 8018
021 3.30 11871
Test 041 2.63 9452
047 1.81 6504
050 3.86 13899
051 0.69 2496
053 2.93 10552
Total 13 32.37 116521
@@ 3.3 Dataset Statistics
Our GraSP dataset comprises thirteen (13) Robot-Assisted Radical Prostatectomy videos captured at a frame rate of 30 frames per second (fps). These surgical videos account for 32.37 hours, with an average duration of 2.49 hours and a standard deviation of 0.77 hours. For benchmarking purposes, we define a two-fold cross-validation setup for training and validation and establish a test set for final performance evaluation. Table 3 presents the data splits and the duration and number of frames (sampled at 1fps) for each video in the GraSP dataset. The original videos of the PSI-AVA dataset with curated annotations make the defined cross-validation set, and the test set comprises the five new videos included in this work.
Figure 1 portrays an overview of our proposed tasks and annotations, and Table 2 presents our sets of semantic categories. The final ground truth annotations of GraSP comprise eleven (11) surgical phase categories, twenty-one (21) surgical steps, seven (7) instrument categories, and fourteen (14) atomic action categories. Therefore, our GraSP dataset proposes diverse classes for each task, integrating local and temporal visual understanding of surgeries. The hierarchical dendrogram of Figure 2 of the Supplementary Material represents the relations between each phase and its nested steps, and it also includes the labels and names of all the categories (IDs) for each task in the GraSP dataset.
@@
Figures 2 and 3 illustrate the distribution of the percentage of annotated frames (counted at 1fps) for each phase and step class, respectively. We also present transition matrices of the annotated categories and visualizations of the temporal progression of phases and steps in Figures 5, 6, 7, and 9 of the Supplementary Material. The frequency distribution of the long-term annotations for a particular class is highly influenced by the occurrence rate of a specific phase or step and the average duration of each occurrence. Notably, these distributions reveal a pronounced long-tail effect, indicating that some phase and step categories are significantly more prevalent than others. For instance, the Idle class is over-represented in the annotations of phases and steps compared to other categories. The transition matrices and temporal visualizations explain this trend, as they demonstrate that most phases and steps frequently transition from or to an Idle state. This tendency is typical and representative of Radical Prostatectomies, as surgeons pause between stages to change instruments, evacuate waste, clean and relocate the endoscope, or rest due to the complexity and duration of the procedures.
Fig. 2: Distribution of the percentage of frames per surgical phase category on each data split. We count the number of frames sampled at 1fps. We organize the graphic in descending order and present correspondence between phase IDs and phase labels in Table 2. The figure is best viewed in color.
Fig. 3: Distribution of the percentage of frames per surgical step category on each data split. We count the number of frames sampled at 1fps. We organize the graphic in descending order and present correspondence between step IDs and step labels in Table 2. The figure is best viewed in color.
@@
Fig. 4: Boxplot distribution of the duration of each phase category. Each boxplot presents the distribution in seconds of the period of all the present temporal windows corresponding to each phase category in our dataset. We clipped this figure to a maximum of 800 seconds for better visualization; the entire figure is presented in the Supplementary Material. The figure is best viewed in color.
@@
Fig. 5: Boxplot distribution of the duration of each step category. Each boxplot presents the distribution in seconds of the period of all the present temporal windows corresponding to each step category in our dataset. We clipped this figure to a maximum of 800 seconds for better visualization; the entire figure is presented in the Supplementary Material. The figure is best viewed in color.
In addition, for Phase and Step annotations, we present the duration distribution of each continuous temporal segment for each phase and step category in seconds. We display the phase duration distribution in Figure 4 and the step duration distribution in Figure 5. These distributions and the visualizations in Figures 7 and 9 of the Supplementary Material underscore these tasks’ granularity and complexity differences. On the one hand, phases represent broader surgery segments, usually occurring once (though interrupted by Idle segments) and covering more general aspects of the procedure. Thus, they constitute a coarser level of understanding and exhibit longer average durations, with some continuous segments lasting over 20 minutes. On the other hand, steps recur in shorter segments repetitively, representing more specific parts of the surgery. Hence, step segments present shorter average durations, with the most extended segments spanning around 11 minutes, indicating a finer-grained level of detail.
Moreover, Figures 4 and 5 denote different intra-class duration variability across all phase and step categories with multiple outliers in most categories. Intuitively, categories with larger average durations exhibit more variability in segment length. For example, step categories such as Identification and Dissection of the iliac vein and artery (1), Cutting and dissection of the external iliac vein’s Lymph Node (2), Ligation of the dorsal venous complex (6), or Cut between the prostate and bladder neck (19) present larger duration times and variabilities as they mainly involve dissection or cutting of blood vessels or tissues and bleeding control, which demand more time, may vary among patients and often require the alternation with other steps. In contrast, categories like Pass suture to the urethra (13), Pass suture to the bladder neck (14), Pull suture (15), or Tie suture (16), show lower durations and variability as they are related to the suturing process which is usually performed within faster and more standardized time windows.
Similar patterns are evident in phase categories where the categories with the highest average duration and variabilities are Left and Right pelvic isolated lymphadenectomy (1 and 2, respectively) and Bladder neck identification and transection (5), which regard tissue cutting and remotion. Conversely, the least durable and variable categories are Prostatic pedicle control (8) and Bladder neck reconstruction (10), mostly related to suturing and clipping procedures but also significantly affected by a consistent interruption by Idle stages. This tendency similarity highlights the complementary relation between phase and step recognition tasks.
In this way, the observed statistical distributions of our long-term annotations, like the long-tail distribution in frame frequency and the significant inter- and intra-class variability in our phases and steps’ duration, present considerable challenges within our benchmark. Nevertheless, it is evident that these distributions accurately represent the realities of surgical scenarios, thus emphasizing the value of utilizing human in vivo data sources and demonstrating the complexities involved in visually understanding surgical scenes.

### v3

@@ 3.1 Data Collection
We recorded 13 surgeries of patients undergoing robotic-assisted laparoscopic Radical Prostatectomy at Fundación Santa Fé de Bogotá, with their explicit consent for data usage. We followed all the data collection protocols with the approval of the ethics committees of all institutions directly involved in this work.
Three highly experienced surgeons performed the surgeries using a da Vinci Si 3000 robot manufactured by Intuitive Inc. (Sunnyvale, California). The equipment included the robot assembled with a high-definition video camera, providing independent viewing for each eye. For recording purposes, we extracted a 2D image signal from the left view of the camera through an HDMI cable connected to a KARL STORZ (AIDA mini WD 100) high-definition video capture device, providing real-time imagery at 720 dots per inch (dpi). We anonymized all the data collected according to our ethical protocols and we manually removed all instrument usage flags or signals in our video frames to avoid biases in model training and evaluation. We provide an example of this frame modification procedure in Figure 1 of the Supplementary Material.
@@
The team placed the Monopolar Energy Scissors (Monopolar Curved Scissors) in port #1 and the Bipolar Energy Forceps (Bipolar Forceps), capable of being fenestrated or configured as Maryland Forceps, in port #2. The third port accommodated a Prograsp Retraction Forceps (Prograsp Forceps). For the reconstructive stages of the surgery, surgeons removed the previous instruments located in ports #1 and #2 and positioned two robotic Needle Holders (Large Needle Drivers). Additionally, the surgical setup designated two ports for the bedside assistant: one port housed a Laparoscopic Suction cannula (Suction Instrument), and the other contained either a Laparoscopic Intestinal Retraction Forceps, a Laparoscopic Needle Holder, a Laparoscopic Scissors designed explicitly for cutting suture material (Laparoscopic Instruments), or a Clip Applier with Plastic Hemostatic clips of the HEM-O-LOCK type.
3.2 Long-Term Tasks Annotation
@@
We categorize phases and steps as long-term reasoning tasks, following the temporal breakdown outlined in [36]. The definitions of phase and step categories and their respective nomenclature adhered to the procedural framework of radical prostatectomy with bilateral extended lymphadenectomy, as validated by the guidelines in [67]. Within this framework, defining surgical phases involved establishing the main stages within the surgical process, each occurring singularly and with minimum repetition. In contrast, surgical steps refer to specific acts undertaken during the surgical process, often occurring multiple times. Hence, the aggregation of multiple steps forms a phase and the length of a phase correlates with the amount of steps.
@@
We then established the sequencing and hierarchy of phase and step categories based on standardized stages of Robot-Assisted Radical Prostatectomy, as routinely practiced at Fundación Santa Fé de Bogotá. Subsequently, two urologists specializing in Robotic-Assisted Surgery reviewed and refined the final category sets, which we portray in Table 2, and their hierarchical structure is visually represented in Figure 2 in the Supplementary Material.
@@
After establishing the labels and hierarchies of the long-term tasks, we conducted an annotation procedure that involved a team of two urologist residents specializing in transperitoneal radical prostatectomy surgeries. Each resident independently annotated all surgical videos, defining specific timestamps for the start and end times of phases and step intervals down to the hour, minute, and second, resulting in two annotations per video. In the case of the eight previously annotated surgeries, these were annotated by the team once more, also resulting in two sets of annotations per surgery. When discrepancies arose between the annotations, the resident annotators collectively reviewed and reconciled differences to establish a consensus on the final annotations.
3.2.1 Long-Term Annotation Consistency Analysis
To evaluate the consistency between our annotators, we analyzed the percentage of frames with differing annotations for each annotator. For phase labels, we observed differing annotations in 21% of frames overall, with an average of 20% differing frames per video and a standard deviation of 9.33%. Step labels exhibited a higher rate, with differing annotations in 37% of frames overall, averaging 36% per video and a standard deviation of 11.33%. These statistics indicate a fair annotator agreement, as the majority of frames were consistently annotated, thus supporting the reliability of our dataset. The increased percentage of differing step annotations reflects their finer granularity and shorter temporal segments, making them more prone to interpretative differences and temporal misalignments. Furthermore, longer videos showed higher rates of differing annotations, likely due to the increased potential for misalignment over extended segments. Despite these challenges, the annotation team’s conciliation process ensured the final dataset’s accuracy and quality for model training and evaluation.
3.3 Short-Term Tasks Annotation
We define Surgical Instrument Segmentation and Atomic Action Detection as short-term, frame-wise, and spatial localization tasks. To ensure balanced category representation while optimizing computational and annotation resources, we only annotated keyframes sampled at 35-second intervals throughout the surgical videos. Initially, the annotations for Instrument Segmentation and Atomic Action Detection were performed independently by separate teams but were later unified into a single dataset, as described in the following subsections. Throughout the process, all annotators had access to 3-second clips centered on each keyframe and full videos for their assigned keyframes to support accurate short-term task annotation. We provide a visual diagram of our short-term annotation process in Figure 3 of the Supplementary Material.
@@ 3.3.1 Surgical Instrument Segmentation
The annotation process for instrument segmentation followed a structured, multi-step approach to ensure high-quality and consistent annotations across the dataset. First, we defined surgical instrument classes based on the da Vinci Robot user manual ([37, 38]) and the instrument setup described in Section 3.1, with the final categories detailed in Table 2. Second, we trained a team of 17 annotators using a standardized and comprehensive tutorial on instrument delineation and classification, followed by multiple annotation tests on challenging frames (e.g., Figure 5 of the Supplementary Material) where annotators had to consistently produce high-quality segmentation masks.
Subsequently, each annotator was randomly assigned a similar number of keyframes to ensure balanced workloads. All annotators were provided with standardized annotation resources, including the open-source platforms Label-Studio ([82]) and Toronto Annotation Suite ([42]), instrument bounding boxes from surgeries annotated in [84], and mask predictions generated by the MATIS model ([5]) pre-trained on Endovis 2017 ([3]) and 2018 ([2]). Annotators were free to utilize any of these resources to delineate and label instrument instances on each keyframe, choosing to either draw and tag masks from scratch, refine the bounding boxes provided by [84] while verifying instrument categories and missing instances, or enhance and label MATIS-generated predictions. Notably, annotators heavily relied on the 3-second clips centered on each keyframe for precision, with most masks being manually drawn with the assistance of each platform’s annotation tools.
After the initial round of segmentation annotations, four surgical data science experts conducted a thorough review to identify errors in classification, delineation, or missed instances. The surgical data scientist then meticulously documented these issues for further correction. We found errors in 9% of final instances, with 8.2% attributed to delineation issues such as imprecise boundaries, inclusion of blood or tissue within the mask, or failure to exclude openings in instrument tips. Misclassifications accounted for 0.3% of instances, while 0.6% involved missing instances. These flagged instances were returned to the annotators for a second round of re-annotation. Following this re-annotation phase, the surgical data science team performed a final validation review of the entire dataset, identifying and personally correcting minor delineation errors in 2% of the instances. While delineation errors are expected due to the inherent complexity of segmentation tasks, our iterative, multi-staged annotation protocol effectively minimized all error types, ensuring high-quality and reliable annotations.
@@
Table 3: Distribution of the number of frames on each defined dataset split. The Table presents the duration in seconds and the number of frames sampled at 1fps and 30fps of each collected Case in our dataset. Each Case corresponds to an entire surgery performed. We also show our predefined dataset partition into Fold 1 and Fold 2 for two-fold cross-validation and a Test set for final model testing.
Set Case No. Duration (h) Frames 1fps Frames 30fps
Fold 1 001 3.05 10972 329159
004 2.36 8490 254678
014 2.77 9973 299118
015 2.48 8969 269031
Fold 2 002 2.62 9443 283261
003 1.63 5882 176453
007 2.23 8018 240522
021 3.30 11871 356097
Test 041 2.63 9452 283564
047 1.81 6504 195131
050 3.86 13899 416973
051 0.69 2496 74876
053 2.93 10552 316552
Total 13 32.37 116521 3495415
3.3.2 Atomic Action Detection
The annotation process for atomic action detection was adapted from the methodology presented in [34]. Our surgical experts first defined a detailed set of atomic actions representing the most detailed actions performed by instruments during surgery, treating these actions as localized, discrete, and instrument-centric states that could be undertaken simultaneously by each instrument. The final atomic action categories are shown in Table 2. Additionally, the surgical team trained the surgical data scientists to identify each atomic action accurately. Alternatively, the surgical team specified which actions could be performed simultaneously by the same instrument and which types of instruments could perform each action. We further organized this information into theoretical co-occurrence matrices (Figure 13 in the Supplementary Material) that we later used to validate the annotations.
Furthermore, the annotation process was conducted per frame and per instance, meaning that each instrument in every keyframe was labeled independently. The surgical data scientists used the Toronto Annotation Suite ([42]) to draw bounding boxes around each instrument instance in the keyframes. To determine the atomic actions performed by each instrument, annotators relied on the 3-second video clips centered on each keyframe, identifying the instrument’s activity at the exact keyframe moment (1.5 seconds into the clip). Thus, each bounding box per keyframe was annotated with up to three atomic action categories, as each instrument instance can perform a maximum of three atomic actions simultaneously.
To ensure annotation robustness and reduce biases, each instrument bounding box was independently labeled three times by different surgical data scientists, resulting in three sets of atomic action labels per instance. The final set of labels for each instance was determined through a consensus approach by retaining only the actions present in at least two of the three annotation sets.
For 2.18% of instances, the overlap between the three annotation sets was notably low (intersection-over-union of two sets over three below 0.3). These cases were carefully reviewed by the surgical data science team, who collaborated to establish the final combination of atomic action labels. Additionally, 5.4% of instances did not align with the established theoretical co-occurrence matrixes. These discrepancies mostly arose when specific actions that could not logically occur in isolation were left as standalone labels after the consensus process. To solve these inconsistencies, the surgical data science team thoroughly reviewed the three original sets of atomic action labels for these instances to understand the source of the inconsistency and defined the final set of atomic action labels. Expert surgeons assisted in all the atomic action annotation review processes, ensuring medical consistency on final annotations.
Fig. 3: Distribution of the percentage of frames per surgical phase category on each data split. We organize the graphic in descending order and present correspondence between phase IDs and phase labels in Table 2. The figure is best viewed in color.
3.3.3 Instrument and Atomic Actions Annotation Alignment
We aligned the atomic action annotations with the instrument segmentation annotations to produce a unified dataset. We achieved this alignment through a bipartite matching per keyframe, where we maximized the Intersection over Union (IoU) between the atomic action bounding boxes and the bounding boxes derived from the boundaries of the segmentation masks. This matching process also served as an additional validation step as it enabled a direct comparison between the instrument masks and the bounding boxes annotated by different teams to identify discrepancies in the number and size of annotated instances. During this process, we identified 2% of the final instances missing from the atomic action annotations and 1.2% from the segmentation annotations, which were mostly highly occluded instances. These instances were re-annotated by the surgical data scientists for both tasks to ensure completeness. Ultimately, we kept the bounding boxes derived from segmentation as the final boxes for atomic action annotations. These boxes adhered to the segmentation masks, making them more precise than those drawn during the atomic action annotation process. This procedure resulted in a cohesive dataset with unified and consistent annotations for instrument instance segmentation and atomic action detection.
Fig. 4: Distribution of the percentage of frames per surgical step category on each data split. We organize the graphic in descending order and present correspondence between step IDs and step labels in Table 2. The figure is best viewed in color.
@@ 3.4 Dataset Statistics
Our GraSP dataset consists of 13 Robot-Assisted Radical Prostatectomy videos, all captured at 30 fps. The total duration of these surgical videos is 32.37 hours, with an average video length of 2.49 hours and a standard deviation of 0.77 hours. For benchmarking purposes, we establish a two-fold cross-validation setup for training and validation and a separate test set for final performance evaluation. The original videos of the PSI-AVA dataset with curated annotations make the cross-validation set, and the test set comprises the five new videos included in this work. Table 3 presents the dataset splits, including the duration and the total number of frames available, sampled at 30 fps and 1 fps. While the frames sampled at 30 fps are available for temporal window creation, the 1 fps frames serve as the primary frames for model evaluation and statistical analysis.
Additionally, Figure 1 and Figure 4 in the Supplementary Material offer an overview of our proposed tasks and annotations. Table 2 lists the semantic categories we have established. Similarly, the hierarchical dendrogram in Figure 2 of the Supplementary Material illustrates the relationships between each surgical phase and its corresponding nested steps.
@@
Figures 3 and 4 illustrate the distribution of the percentage of annotated frames for each phase and step class, respectively. We also present the number of frames per phase and step category in Tables 1 and 2 of the Supplementary Material, along with transition matrices of the annotated categories and visualizations of the temporal progression of phases and steps in Figures 7, 8, 9, and 11 of the Supplementary Material. Notably, these distributions exhibit a pronounced long-tail effect, with the Idle class over-represented, as most phases and steps frequently transition from or to an Idle state due to surgeons repeatedly pausing, resting, or adjusting the surgical setup between stages.
In addition, we present the duration distribution, in seconds, of each continuous temporal segment for each phase and step category in Figures 5 and 6. These distributions and the visualizations in Figures 9 and 11 of the Supplementary Material underscore the granularity difference between phases and steps: phases represent broader surgery segments with longer average durations, with some segments lasting over 20 minutes. Conversely, steps recur in shorter and repetitive segments, with the most extended segments spanning around 10 minutes. Figures 5 and 6 also denote different intra-class duration variability across all phase and step categories with multiple outliers. Intuitively, categories involving tissue dissection, vessel cutting, or bleeding control (e.g., Phases 1-3, Steps 1-3,5 & 6) often present longer durations and variability as they demand more time, may vary among patients, and often require alternation between steps. In contrast, categories related to the suturing process (e.g., Phase 8, Steps 13-16) show shorter durations and variability as these are usually performed within faster and more standardized time windows. These statistical distributions of our long-term annotations demonstrate substantial challenges within our benchmark and closely reflect the realities of surgical scenarios.
Furthermore, Figure 7 and Table 3 (Supplementary Material) display the frequency distribution of instrument instances per class, also demonstrating a natural long-tail effect. This distribution is due to robotic instruments like BF, and MCS being more actively used owing to their higher versatility and action capacity than LND or PF, which are more specialized. In contrast, laparoscopic instruments like CA or LI are additional to the da Vinci Surgical System (dVSS) and are employed for specific purposes during particular stages of the surgery. Moreover, Figure 8 and Table 4 (Supplementary Material) illustrate the distribution of the number of instrument instances per video keyframe. We observe that each keyframe contains 1 to 5 instrument instances (2.23 average of instrument instances per frame), with approximately 80% of frames containing between 2 and 3 instances. These statistics are characteristic of the dVSS, typically involving two robotic instruments operated by the surgeon with occasional stationary robotic tools or additional non-robotic tools handled by an assistant that increase the complexity of surgical scenes.
@@
Fig. 5: Boxplot distribution of the duration of each phase category. Each boxplot presents the distribution in seconds of the period of all the present temporal windows corresponding to each phase category in our dataset. We clipped this figure to a maximum of 800 seconds for better visualization; the entire figure is presented in Figure 10 of the Supplementary Material. The figure is best viewed in color.
@@
Fig. 6: Boxplot distribution of the duration of each step category. Each boxplot presents the distribution in seconds of the period of all the present temporal windows corresponding to each step category in our dataset. We clipped this figure to a maximum of 800 seconds for better visualization; the entire figure is presented in Figure 12 of the Supplementary Material. The figure is best viewed in color.